At the center of a sphere of radius 0.5m a point charge of 2μC is placed. What electric flux is passing through the sphere?

1
Expert's answer
2022-03-09T10:32:49-0500

The electric flux through the surface


"\Phi=\oint EdA=\frac{q}{\epsilon_0}""\Phi=\frac{2.0*10^{−6}}{8.85*10^{-12}}=2.3*10^{5}\:\rm N\cdot m^2/C"